At the instant shown, the arm ?? has an angular velocity of ?⃗ = −4 ?̂ ??? ? , and no angular acceleration.

1€ 4 m verur - - - - - - 3 m 1 m sw=-4 Å RAD LAL a=o

a. Determine the velocity of the block C and the angular velocity of the arm BC.

b. Determine the angular acceleration of the arm ?? and the acceleration of the block ?.
Answer #1

Arm AB VB = VA + WAB XY AB = 0 - 4k² x (H) VB = 4 i mis 2 AB = AA + LABK X8AB - (WAB) VAB + LABK X SAB - 10 Op = 0 +0 =(-4)